OpenStudy (anonymous):

given the implicit relation (x^2)-2xy+(y^3)=0, used partial derivative to calculate dy/dx

OpenStudy (anonymous):

\[dy/dx=-F _{x}/F _{y}\]

OpenStudy (anonymous):

confuse

OpenStudy (anonymous):

\[dy/dx=-(2x-2y)/(-2x+3y ^{2})\]

OpenStudy (anonymous):

how did u get get rid of the 0

OpenStudy (anonymous):

to find this derivative you have to use implicit diferentiation , or to say it other way, make derivative of x considering y a function of x.\[F _{x}+F _{y}*y'=0\] where y' =dy/dx now pass The Fx to left and devide everything by Fy and you got it

OpenStudy (anonymous):

can u show it step by step

OpenStudy (anonymous):

this is step by step

OpenStudy (anonymous):

with (x^2)-2xy+(y^3)=0

OpenStudy (anonymous):

\[F _{x}=2x-2y\] \[F _{y}=-2x+3y ^{2}\] later just use the formula: \[dy/dx=-F _{x}/F _{y}=-(2x-2y)/(-2x+3y ^{2})\]

OpenStudy (anonymous):

thanks